# Concentration and fluctuations of sine–Gordon measure around topological multi-soliton manifold

### v1

@@ 5. Gaussian measures associated with Schrödinger operators
In this section, we study the linearized operator around the multi-soliton configuration. For well-separated centers $\xi_{1},\dots,\xi_{k}$ satisfying
$\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\to\infty$
as $\varepsilon\to 0$, the Hessian of the energy at $m_{\xi_{1},\dots,\xi_{k}}=\sum_{j=1}^{k}m_{\xi_{j}}$ is
@@
$=-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\textup{sech}^{2}(\cdot-\xi_{j})+O(e^{-c\min_{i\neq j}|\xi_{i}-\xi_{j}|}),$ (5.1)
@@
$\mathcal{L}_{\xi}=-\partial_{x}^{2}+\cos(m_{\xi})=-\partial_{x}^{2}+1-2\text{sech}^{2}(\cdot-\xi),$ (5.2)
@@
    $\langle\mathcal{L}_{\xi}v,v\rangle\geq\lambda_{0}\|v\|_{H^{1}}^{2}.$ (5.4)
@@
$\mathcal{L}_{\xi_{1},\dots,\xi_{k}}:=-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j}).$
@@
where we used $\mathcal{L}_{\xi_{j}}(\partial_{\xi_{j}}m_{\xi_{j}})=0$ from (5.3). Hence, the linearized operator $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}$ satisfies
@@
$\geq\sum_{j}\langle\mathcal{L}_{\xi_{1},\dots,\xi_{k}}(\chi_{j}v),\chi_{j}v\rangle_{L^{2}(\mathbb{R})}-Cd_{\varepsilon}^{2}\|v\|_{L^{2}}^{2},$ (5.5)
@@
where $\mathcal{L}_{\xi_{j}}=-\partial_{x}^{2}+1-2\text{sech}^{2}(\cdot-\xi_{j})$. This implies that
$\langle\mathcal{L}_{\xi_{1},\dots,\xi_{k}}(\chi_{j}v),\chi_{j}v\rangle_{L^{2}(\mathbb{R})}\geq\langle\mathcal{L}_{\xi_{j}}(\chi_{j}v),\chi_{j}v\rangle_{L^{2}(\mathbb{R})}-Ce^{-cd_{\varepsilon}}\|\chi_{j}v\|_{L^{2}}^{2}.$ (5.6)
@@
$|\langle\chi_{j}v,\partial_{\xi_{j}}m_{\xi_{j}}\rangle_{L^{2}(\mathbb{R})}|\lesssim\sum_{i:i\neq j}e^{-cd_{\varepsilon}}\|\chi_{i}v\|_{L^{2}}^{2}.$ (5.7)
@@
$\langle\mathcal{L}_{\xi_{j}}(\chi_{j}v),\chi_{j}v\rangle_{L^{2}(\mathbb{R})}\geq\frac{\lambda_{0}}{2}\|\chi_{j}v\|_{H^{1}}^{2}-Ce^{-cd_{\varepsilon}}\sum_{i:i\neq j}\|\chi_{i}v\|_{L^{2}}^{2}$ (5.8)
@@
From now on, we study the operator
$\mathcal{L}_{\xi_{1},\dots,\xi_{k}}=-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$
on the finite interval $L_{\varepsilon},L_{\varepsilon}]$ with Dirichlet boundary conditions.
@@
$\langle\mathcal{L}_{\xi_{1}^{n},\dots,\xi_{k}^{n}}\widetilde{v}_{n},\widetilde{v}_{n}\rangle_{L^{2}(\mathbb{R})}=\langle\mathcal{L}_{\xi_{1}^{n},\dots,\xi_{k}^{n}}v_{n},v_{n}\rangle_{L^{2}([-L_{n},L_{n}])}\leq\frac{1}{n}\|v_{n}\|^{2}_{L^{2}([-L_{n},L_{n}])}=\frac{1}{n}\|\widetilde{v}_{n}\|^{2}_{L^{2}(\mathbb{R})},$ (5.9)
@@
$\langle\mathcal{L}_{\xi_{1}^{n},\dots,\xi_{k}^{n}}\widetilde{v}_{n},\widetilde{v}_{n}\rangle_{L^{2}(\mathbb{R})}\geq\lambda_{0}\|\widetilde{v}_{n}\|_{L^{2}(\mathbb{R})}^{2}.$ (5.10)
@@
$C_{\xi_{1},\dots,\xi_{k}}=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})\big)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}},$ (5.11)
@@
$\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}=\operatorname{Id}-\sum_{j=1}^{k}\langle\cdot,t_{j}\rangle t_{j}=\operatorname{Id}-\sum_{j=1}^{k}\mathbf{P}_{j}$ (5.12)
@@
$\mathbf{P}_{j}=\langle\cdot,t_{j}\rangle t_{j}.$
@@
The operator
$-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$
considered on the finite interval $L_{\varepsilon},L_{\varepsilon}]$, is a self-adjoint Sturm–Liouville operator since $2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$ is a Schwartz function. Hence, the operator has purely discrete spectrum $\{\lambda_{n}\}_{n}$ with $\lambda_{n}\sim|n|^{2}$ as $n\to\infty$. In particular, if the operator is strictly positive, its inverse belongs to the trace class. By Lemma 5.3, the projected operator is strictly positive, hence its inverse on that subspace is trace class. Therefore, the corresponding Gaussian measure is well defined.
@@
Let $Q\in\mathbb{Z}$ with $|Q|=k$ and $d_{\varepsilon}=|\log(\varepsilon\log\frac{1}{\varepsilon})|$. Then we have
@@
uniformly in $\xi_{1},\dots,\xi_{k}$ satifying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}$, where $Z_{0}$ is the partition function of the Gaussian measure associated with the single-well operator $\mathcal{L}_{0}=-\partial_{x}^{2}+1-\textup{sech}^{2}(x)$.
@@
$1\leq\lambda_{1}<\lambda_{2}<\dots<\lambda_{\ell}<\dots\to\infty.$ (5.13)
@@
$|\phi_{\ell}(x)|+|\partial_{x}\phi_{\ell}(x)|\lesssim e^{-c|x|}.$ (5.14)
For each $\xi_{j}$, $j=1,\dots,k$, we define
$\phi_{\ell,j}(x):=\phi_{\ell}(x-\xi_{j}).$
We can easily verify that $\phi_{\ell,j}$ is an eigenfunction of $\mathcal{L}_{\xi_{j}}=-\partial_{x}^{2}+1-\textup{sech}^{2}(\cdot-\xi_{j})$ with the same eigenvalue $\lambda_{\ell}$. Furthermore the translated eigenfunctions $\phi_{\ell,j}$ act as approximate eigenfunctions for $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}=-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\textup{sech}^{2}(\cdot-\xi_{j})$ as follows
@@
$\lambda_{\ell,j}=\lambda_{\ell,j}(\xi_{1},\dots,\xi_{k})=\lambda_{\ell}+O(e^{-cd_{\varepsilon}}),\quad j=1,\dots,k.$ (5.15)
@@
uniformly in $\xi_{1},\dots,\xi_{k}$ satifying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}$, where $\mathbf{P}_{N}$ is the finite-dimensional projection onto modes with $|\ell|\leq N$.
6. Concentration around the multi-soliton manifold without collisions
In this section, we prove Theorem 1.2.
6.1. Free energy
Before proving Theorem 1.2, we study in this subsection the behavior of the free energy as $\varepsilon\to 0$.
@@ 6.1. Free energy
$\lim_{\varepsilon\to 0}\varepsilon\log Z_{\varepsilon}^{Q}=-\inf_{\phi\in\mathcal{C}_{Q}}E(\phi),$
where $\mathcal{C}_{Q}$ is the topological sector in (1.3) and $Z^{Q}_{\varepsilon}$ is the partition function of the Gibbs measure $\rho^{Q}_{\varepsilon}$ defined in (1.5).
Remark 6.2.
Based on Lemma 2.2 (especially (2.8)), a naive calculation would restrict the interval size to $L_{\varepsilon}=\varepsilon^{-\frac{1}{3}+\eta}$. However, by exploiting the specific structure of the sine–Gordon model, we are able to extend the interval up to $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$. See (6.6), (6.7), and (6.8).
@@
By using the Boué–Dupuis formula, Lemma 2.1, we write
$\log Z^{Q}_{\varepsilon}$ $=\log\mathbb{E}_{\mu_{\varepsilon}^{Q}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\bigg]$
$=\sup_{\theta\in\mathbb{H}_{a}}\mathbb{E}\bigg[-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\big(1-\cos(\ell^{Q}+\sqrt{\varepsilon}Y+\sqrt{\varepsilon}\Theta)\big)dx-\frac{1}{2}\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt\bigg],$
where we used the notation $Y:=Y(1)$ and $\Theta:=\Theta(1)$ for convenience.
We first derive an upper bound on the free energy. By performing the change of variables $\sqrt{\varepsilon}\theta\to\theta$, which does not affect the variational problem, we can rewrite
$\varepsilon\log Z^{Q}_{\varepsilon}$ $=\sup_{\theta\in\mathbb{H}_{a}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+\ell^{Q}+\sqrt{\varepsilon}\Theta))dx-\frac{\varepsilon}{2}\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt\bigg]$
$=\sup_{\theta\in\mathbb{H}_{a}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta))dx-\frac{1}{2}\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt\bigg]$
$\leq\sup_{\Theta\in\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta))dx-\frac{1}{2}\|\partial_{x}\Theta\|_{L^{2}}^{2}\bigg],$ (6.1)
@@
$\hphantom{XXXXXXXXXXXX}-\frac{1}{2}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\partial_{x}W\cdot\partial_{x}\ell^{Q}dx\bigg].$ (6.3)
@@
$\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\partial_{x}W\cdot\partial_{x}\ell^{Q}dx\bigg|\leq\frac{\delta}{2}\|\partial_{x}W\|_{L^{2}}^{2}+C_{\delta}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2},$ (6.4)
@@
$\varepsilon\log Z_{\varepsilon}^{Q}$ $\leq\sup_{W\in\ell^{Q}+\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+W))dx-\frac{1-\delta}{2}\|\partial_{x}W\|_{L^{2}}^{2}\bigg]+C_{\delta}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}.$ (6.5)
@@
$\cos(\sqrt{\varepsilon}Y+W)=\cos(W)-\sin(W)\cdot(\sqrt{\varepsilon}Y)+O\big((\sqrt{\varepsilon}Y)^{2}\big).$ (6.6)
Note that $\sin^{2}(W)=(1-\cos(W))(1+\cos(W))\leq 2(1-\cos(W))$. Hence,
$|\sin(W)\cdot(\sqrt{\varepsilon}Y)|\leq\kappa(1-\cos(W))+C_{\kappa}(\sqrt{\varepsilon}Y)^{2}$ (6.7)
for an arbitrarily small $\kappa>0$ and a large constant $C_{\kappa}$. By plugging (6.6) into (6.5) and using (6.7), we obtain
@@
$\leq\sup_{W\in\ell^{Q}+\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\kappa)(1-\cos W)dx-\frac{1-\delta}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}W|^{2}dx\bigg]+C_{\kappa}\varepsilon L_{\varepsilon}^{2}+C_{\delta}\frac{(\pi Q)^{2}}{L_{\varepsilon}}.$ (6.8)
@@
$\limsup_{\varepsilon\to 0}\varepsilon\log Z_{\varepsilon}^{Q}\leq-\inf_{\phi\in\mathcal{C}_{Q}}E(\phi).$
@@
$\varepsilon\log Z_{\varepsilon}^{Q}=\sup_{\theta\in\mathbb{H}_{a}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta))dx-\frac{1}{2}\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt\bigg].$ (6.9)
@@
$\theta^{0}(t)=\eta^{-1}\cdot\mathbf{1}_{\{t>1-\eta\}}(t)(-\partial_{x}^{2})^{\frac{1}{2}}(-\sqrt{\varepsilon}Y_{N}(1-\eta)-\ell^{Q}+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}),$ (6.10)
where $\eta>0$, $\xi_{1},\dots,\xi_{k}\in[-L_{\varepsilon},L_{\varepsilon}]$, and
$Y_{N}(1-\eta)=\sum_{1\leq n\leq N}\frac{B_{n}(1-\eta)}{\sqrt{\lambda_{n}}}e_{n}(x),$

### v2

@@ 5. Gaussian measures associated with Schrödinger operators
In this section, we study the linearized operator around the multi-soliton configuration. For well-separated centers $\xi_{1},\dots,\xi_{k}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\to\infty$, the Hessian of the energy at $m_{\xi_{1},\dots,\xi_{k}}=\sum_{j=1}^{k}m_{\xi_{j}}$ is
@@
(5.1) $=-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\textup{sech}^{2}(\cdot-\xi_{j})+O(e^{-c\min_{i\neq j}|\xi_{i}-\xi_{j}|}),$
@@
(5.2) $\mathcal{L}_{\xi}=-\partial_{x}^{2}+\cos(m_{\xi})=-\partial_{x}^{2}+1-2\text{sech}^{2}(\cdot-\xi),$
@@
    (5.4) $\langle\mathcal{L}_{\xi}v,v\rangle\geq\lambda_{0}\|v\|_{H^{1}}^{2}.$
@@
(5.5) $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}:=-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j}).$
@@
where we used $\mathcal{L}_{\xi_{j}}(\partial_{\xi_{j}}m_{\xi_{j}})=0$ from (5.3). Hence, the linearized operator $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}$ in (5.5) satisfies
@@
(5.6) $\geq\sum_{j}\langle\mathcal{L}_{\xi_{1},\dots,\xi_{k}}(\chi_{j}v),\chi_{j}v\rangle_{L^{2}(\mathbb{R})}-Cd_{\varepsilon}^{2}\|v\|_{L^{2}}^{2},$
@@
where $\mathcal{L}_{\xi_{j}}$ is defined as in (5.2). This implies that
(5.7) $\langle\mathcal{L}_{\xi_{1},\dots,\xi_{k}}(\chi_{j}v),\chi_{j}v\rangle_{L^{2}(\mathbb{R})}\geq\langle\mathcal{L}_{\xi_{j}}(\chi_{j}v),\chi_{j}v\rangle_{L^{2}(\mathbb{R})}-Ce^{-cd_{\varepsilon}}\|\chi_{j}v\|_{L^{2}}^{2}.$
@@
(5.8) $|\langle\chi_{j}v,\partial_{\xi_{j}}m_{\xi_{j}}\rangle_{L^{2}(\mathbb{R})}|\lesssim\sum_{i:i\neq j}e^{-cd_{\varepsilon}}\|\chi_{i}v\|_{L^{2}}^{2}.$
@@
(5.9) $\langle\mathcal{L}_{\xi_{j}}(\chi_{j}v),\chi_{j}v\rangle_{L^{2}(\mathbb{R})}\geq\frac{\lambda_{0}}{2}\|\chi_{j}v\|_{H^{1}}^{2}-Ce^{-cd_{\varepsilon}}\sum_{i:i\neq j}\|\chi_{i}v\|_{L^{2}}^{2}$
@@
From now on, we study the operator $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}$, defined in (5.5), on the finite interval $L_{\varepsilon},L_{\varepsilon}]$ with Dirichlet boundary conditions.
@@
(5.10) $\langle\mathcal{L}_{\xi_{1}^{n},\dots,\xi_{k}^{n}}\widetilde{v}_{n},\widetilde{v}_{n}\rangle_{L^{2}(\mathbb{R})}=\langle\mathcal{L}_{\xi_{1}^{n},\dots,\xi_{k}^{n}}v_{n},v_{n}\rangle_{L^{2}([-L_{n},L_{n}])}\leq\frac{1}{n}\|v_{n}\|^{2}_{L^{2}([-L_{n},L_{n}])}=\frac{1}{n}\|\widetilde{v}_{n}\|^{2}_{L^{2}(\mathbb{R})},$
@@
(5.11) $\langle\mathcal{L}_{\xi_{1}^{n},\dots,\xi_{k}^{n}}\widetilde{v}_{n},\widetilde{v}_{n}\rangle_{L^{2}(\mathbb{R})}\geq\lambda_{0}\|\widetilde{v}_{n}\|_{L^{2}(\mathbb{R})}^{2}.$
@@
(5.12) $C_{\xi_{1},\dots,\xi_{k}}=\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}\big(-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})\big)^{-1}\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}},$
@@
(5.13) $\mathbf{P}_{V_{\xi_{1},\dots,\xi_{k}}}=\operatorname{Id}-\sum_{j=1}^{k}\langle\cdot,t_{j}\rangle t_{j}=\operatorname{Id}-\sum_{j=1}^{k}\mathbf{P}_{j},$
@@
(5.14) $\mathbf{P}_{j}=\langle\cdot,t_{j}\rangle t_{j}.$
@@
The operator $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}=-\partial_{x}^{2}+1-2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$, considered on the finite interval $L_{\varepsilon},L_{\varepsilon}]$, is a self-adjoint Sturm–Liouville operator since $2\sum_{j=1}^{k}\text{sech}^{2}(\cdot-\xi_{j})$ is a Schwartz function. Hence, the operator has purely discrete spectrum $\{\lambda_{n}\}_{n}$ with $\lambda_{n}\sim|n|^{2}$ as $n\to\infty$. In particular, if the operator is strictly positive, its inverse belongs to the trace class. By Lemma 5.3, the projected operator is strictly positive, hence its inverse on that subspace is trace class. Therefore, the corresponding Gaussian measure is well defined.
@@
Let $Q\in\mathbb{Z}$ with $|Q|=k$. Then we have
@@
uniformly in $\xi_{1},\dots,\xi_{k}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}\to\infty$ as $\varepsilon\to 0$, where $Z_{0}$ is the partition function of the Gaussian measure associated with the single-well operator $\mathcal{L}_{0}=-\partial_{x}^{2}+1-\textup{sech}^{2}(x)$.
@@
(5.15) $1\leq\lambda_{1}<\lambda_{2}<\dots<\lambda_{\ell}<\dots\to\infty.$
@@
(5.16) $|\phi_{\ell}(x)|+|\partial_{x}\phi_{\ell}(x)|\lesssim e^{-c|x|}.$
For each $\xi_{j}$, $j=1,\dots,k$, we define $\phi_{\ell,j}(x):=\phi_{\ell}(x-\xi_{j})$. We can easily verify that $\phi_{\ell,j}$ is an eigenfunction of $\mathcal{L}_{\xi_{j}}$, defined in (5.2), with the same eigenvalue $\lambda_{\ell}$. Furthermore the translated eigenfunctions $\phi_{\ell,j}$ act as approximate eigenfunctions for $\mathcal{L}_{\xi_{1},\dots,\xi_{k}}$, defined in (5.5), as follows
@@
(5.17) $\lambda_{\ell,j}=\lambda_{\ell,j}(\xi_{1},\dots,\xi_{k})=\lambda_{\ell}+O(e^{-cd_{\varepsilon}}),\quad j=1,\dots,k.$
@@
uniformly in $\xi_{1},\dots,\xi_{k}$ satisfying $\min_{i\neq j}|\xi_{i}-\xi_{j}|\geq d_{\varepsilon}$, where $\mathbf{P}_{N}$ is the finite-dimensional projection onto the modes with $|\ell|\leq N$.
6. Proof of Theorem 1.2 (concentration around non-collision manifold)
In this section, we prove Theorem 1.2. Note that
$\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)=\frac{Z_{\varepsilon}^{Q}\Big[\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}}\Big]}{Z_{\varepsilon}^{Q}[1]}$
where
$Z_{\varepsilon}^{Q}[g]:=\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}g\bigg].$
This implies
(6.1) $\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)=\varepsilon\log Z_{\varepsilon}^{Q}\Big[\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}}\Big]-\varepsilon\log Z_{\varepsilon}^{Q}[1].$
Similarly,
(6.2) $\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta\big)=\varepsilon\log Z_{\varepsilon}^{Q}\Big[\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta\}}\Big]-\varepsilon\log Z_{\varepsilon}^{Q}[1].$
Below we prove the behavior of the free energy $\varepsilon\log Z_{\varepsilon}^{Q}[1]$ as $\varepsilon\to 0$, and also obtain upper and lower bounds on $\varepsilon\log Z_{\varepsilon}^{Q}[g]$, with $g$ being the respective indicator functions in (6.1) and (6.2).
For the free energy, by Lemma 2.1, we write
$\varepsilon\log Z^{Q}_{\varepsilon}[1]$ $=\sup_{\theta\in\mathbb{H}_{a}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\big(1-\cos(\ell^{Q}+\sqrt{\varepsilon}Y+\sqrt{\varepsilon}\Theta)\big)dx-\frac{\varepsilon}{2}\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt\bigg]$
(6.3) $=\sup_{\theta\in\mathbb{H}_{a}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta))dx-\frac{1}{2}\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt\bigg],$
where we used the notation $Y:=Y(1)$ and $\Theta:=\Theta(1)$ for convenience, and in the second line we perform the change of variables $\sqrt{\varepsilon}\theta\to\theta$ which does not affect the variational problem.
Below we prove an upper bound on $\varepsilon\log Z_{\varepsilon}^{Q}[1]$ in terms of the minimal energy (see (6.13)). To this end, we suitably expand the potential in the fluctuation $Y$ as in (6.10). However, in a naive expansion of the potential, one would obtain, at the first order
(6.4) $\mathbb{E}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}Y|dx\sim\varepsilon^{\frac{1}{2}}L_{\varepsilon}^{\frac{3}{2}}$
in view of (2.8) in Lemma 2.2. So for the fluctuation (6.4) to vanish in order to get a bound merely in terms of the minimal energy, one would eventually have to restrict the interval size to $L_{\varepsilon}=\varepsilon^{-\frac{1}{3}+\eta}$, which is much unnatural and non-optimal. For the upper bounds on $\varepsilon\log Z_{\varepsilon}^{Q}[g]$ with $g$ being the respective indicator functions, we would run into the same problem.
However, below we show that by exploiting a simple but crucial specific structure of the sine–Gordon model, we are actually able to extend the interval up to $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+\eta}$. See (6.10), (6.11), and (6.12).
@@
$\lim_{\varepsilon\to 0}\varepsilon\log Z_{\varepsilon}^{Q}[1]=-\inf_{\phi\in\mathcal{C}_{Q}}E(\phi),$
where $\mathcal{C}_{Q}$ is the topological sector in (1.3)
@@
In this proof we write $Z_{\varepsilon}^{Q}=Z_{\varepsilon}^{Q}[1]$. We first derive an upper bound. By (6.3) and (2.9)
(6.5) $\varepsilon\log Z^{Q}_{\varepsilon}$ $\leq\sup_{\Theta\in\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta))dx-\frac{1}{2}\|\partial_{x}\Theta\|_{L^{2}}^{2}\bigg],$
@@
(6.7) $\hphantom{XXXXXXXXXXXX}-\frac{1}{2}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\partial_{x}W\cdot\partial_{x}\ell^{Q}dx\bigg].$
@@
(6.8) $\bigg|\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\partial_{x}W\cdot\partial_{x}\ell^{Q}dx\bigg|\leq\frac{\delta}{2}\|\partial_{x}W\|_{L^{2}}^{2}+C_{\delta}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2},$
@@
(6.9) $\varepsilon\log Z_{\varepsilon}^{Q}$ $\leq\sup_{W\in\ell^{Q}+\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+W))dx-\frac{1-\delta}{2}\|\partial_{x}W\|_{L^{2}}^{2}\bigg]+C_{\delta}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}.$
@@
(6.10) $\cos(\sqrt{\varepsilon}Y+W)=\cos(W)-\sin(W)\cdot(\sqrt{\varepsilon}Y)+O\big((\sqrt{\varepsilon}Y)^{2}\big).$
Note that
$\sin^{2}(W)=(1-\cos(W))(1+\cos(W))\leq 2(1-\cos(W)).$
Hence,
(6.11) $|\sin(W)\cdot(\sqrt{\varepsilon}Y)|\leq\kappa(1-\cos(W))+C_{\kappa}(\sqrt{\varepsilon}Y)^{2}$
for an arbitrarily small $\kappa>0$ and a large constant $C_{\kappa}$. Note that the first term can be absorbed into the potential. By plugging (6.10) into (6.9) and using (6.11), we obtain
@@
(6.12) $\leq\sup_{W\in\ell^{Q}+\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\kappa)(1-\cos W)dx-\frac{1-\delta}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}W|^{2}dx\bigg]+C_{\kappa}\varepsilon L_{\varepsilon}^{2}+C_{\delta}\frac{(\pi Q)^{2}}{L_{\varepsilon}}.$
@@
(6.13) $\limsup_{\varepsilon\to 0}\varepsilon\log Z_{\varepsilon}^{Q}\leq-\inf_{\phi\in\mathcal{C}_{Q}}E(\phi).$
@@
(6.14) $\varepsilon\log Z_{\varepsilon}^{Q}=\sup_{\theta\in\mathbb{H}_{a}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta))dx-\frac{1}{2}\int_{0}^{1}\|\theta(t)\|_{L^{2}_{x}}^{2}dt\bigg].$
@@
(6.15) $\theta^{0}(t)=\zeta^{-1}\cdot\mathbf{1}_{\{t>1-\zeta\}}(t)(-\partial_{x}^{2})^{\frac{1}{2}}(-\sqrt{\varepsilon}Y_{N}(1-\zeta)-\ell^{Q}+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}),$
where $\zeta>0$, $\xi_{1},\dots,\xi_{k}\in[-L_{\varepsilon},L_{\varepsilon}]$, and
$Y_{N}(1-\zeta)=\sum_{1\leq n\leq N}\frac{B_{n}(1-\zeta)}{\sqrt{\lambda_{n}}}e_{n}(x),$